### v2

@@ 7 Single-commodity instances with multiple origins and destinations
Unfortunately, we cannot handle multiple sources so easily. To see why the situation is different here, consider a setting with two sources $s_{1}$ and $s_{2}$ both connected to some vertex $v$, which is then connected to the sink $t$. Let us assume that the travel time from $s_{1}$ to $v$ is much shorter than the travel time from $s_{2}$ to $v$. In this case the particles which enter the network via $s_{2}$ at time 0 interact with particles which enter the network via $s_{1}$ at a much later time. If we were to contract $s_{1}$ and $s_{2}$ together, no particles would depart from $s_{2}$ at time 00, which is certainly incorrect.
Fortunately, the question of how to handle multiple source has already been discussed by Sering and Skutella [SS18]. They showed that Nash flows over time exist in this setting, and show that their label derivatives satisfy a modified version of the thin flow equations almost everywhere. We cannot quite use their results in a black box fashion, since we would like to simply apply our results for the single-source single-sink case, rather then redoing the proofs with slightly different thin flow equations.
Let $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{m}\}$ be the set of sources (and $t$ the sink, as usual), and let $u_{i}$ be the inflow rate at $s_{i}$. It will be convenient to define the following network $\hat{G}$. Its node set $V(\hat{G})$ consists of $V$ along with an additional “super-source” $s$. Its arc set $E(\hat{G})$ consists of $E$, along with the arcs $E_{s}:=\{ss_{i}:s_{i}\in\mathcal{S}\}$. All arcs in $E$ maintain their transit times and capacities; arc $ss_{i}$ has capacity $u_{i}$ and transit time 00 for each $s_{i}\in\mathcal{S}$. Let $u_{0}\coloneqq\sum_{i=1}^{m}u_{i}$.
We will need to consider the details of the modified thin flow equations of [SS18, Definition 3]. The following is easily be seen to be a slightly rewritten version:
@@
[…]
@@
We remark that $\theta$ can no longer be interpreted as a time, since the first particles departing any source depart at time 00, even though the smallest value of $\theta$ such that $x^{\prime}_{ss_{i}}$ is positive can be nonzero and distinct for different sources. Rather, $\theta$ is simply an index into the set of agents; we refer the reader to [SS18] for more clarity on the intuition.
Call a source $s_{i}$ active at $\theta$ if there is a path in $(V,E^{\prime}_{\theta})$ from $s_{i}$ to $t$. Let $\mathcal{S}_{\theta}$ is the set of active sources at time $\theta$. Now let $\mathcal{Z}_{\theta}$ be the set of nodes reachable from the inactive sources $\mathcal{S}\setminus\mathcal{S}_{\theta}$ along a path of active arcs. Then it is easy to observe that in the modified thin flow equations, an inactive source $s_{i}$ must have $x^{\prime}_{ss_{i}}=\ell^{\prime}_{s_{i}}=0$, since there is no flow through $s_{i}$ supported on $E^{\prime}_{\theta}$ It also then follows that all nodes in $v\in\mathcal{Z}_{\theta}$ satisfy $\ell^{\prime}_{v}=0$. Let $E^{\prime\prime}_{\theta}=E^{\prime}_{\theta}\cup\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$, and let $E^{**}_{\theta}:=E^{*}_{\theta}\cup\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$. Then the following is equivalent to (10):
@@
[…]
Here, we have extended $\rho_{e}$ to $e=ss_{i}$ for $s_{i}\in\mathcal{S}_{\theta}$ in the obvious way $\rho_{e}(\ell^{\prime}_{s},x^{\prime}_{e})=x^{\prime}_{e}/u_{i}$, given that these are all considered resetting arcs. Now define the generalized network $G_{\theta}$ by first removing $\mathcal{Z}_{\theta}$ from $\hat{G}$, and then choosing the set of free arcs to be $E^{\infty}_{\theta}:=\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$. Then (11) is identical to the usual thin flow equations (2) on $G_{\theta}$, along with $\ell^{\prime}_{v}=0$ for $v\in\mathcal{Z}_{\theta}$. $(E(G_{\theta}),E^{\infty}_{\theta})$ needs to be a valid configuration: this requirement provides a definition of a feasible labeling in this setting. We need one further restriction: if $s_{i}$ is in inactive source at time $\theta$, then $\ell_{s_{i}}(\theta)=0$. As usual, we can consider equilibrium trajectories starting from any feasible labeling $l^{\circ}$.
@@
[…]
@@
    For all points of differentiability $\theta$ and $s_{i}\in\mathcal{S}\setminus\mathcal{S}_{\theta}$, $\beta_{s_{i}}^{\prime}(\theta)\leq-\kappa_{\min}$.
@@
    For any $e=vw\in\tilde{P}$, since $e$ has no queue, we have $s^{\prime}_{e}(\theta)=\ell_{v}^{\prime}(\theta)-\ell_{w}^{\prime}(\theta)$ (recall that $\theta$ is a point of differentiability). Define $\kappa_{\min}$ to be the minimum possible value of $\ell^{\prime}_{v}$ for any node $v$ and any thin flow solution $\ell^{\prime}$ in $G_{\theta}$, over all possible valid configurations (this is strictly positive by Lemma 4.4 and since no node of $\mathcal{Z}_{\theta}$ is in $G_{\theta}$). Then
@@
    Note that a source $s_{i}$ is active when $\beta_{s_{i}}(\theta)=0$. It follows directly that the set of active sources $\mathcal{S}_{\theta}$ only increases in $\theta$, since $\beta_{s_{i}}$ is clearly continuous.
    Next, observe that for every inactive source $s_{i}$ the maximal slack on a path to the sink is upper bounded by $\sum_{e\in E}\tau_{e}$. Let $P$ be an $s_{i}$-$t$-path of minimum slack at time 00 and let $\tilde{P}$ be a maximal prefix of $P$ such that no arc of $\tilde{P}$ has a queue. Then (similarly to the above)
@@
By Lemma 7.1, we have that for all $\theta\geq\sum_{e\in E}\tau_{e}/\kappa_{\min}$ all sources are active and remain active. Thus $\mathcal{Z}_{\theta}=\emptyset$, and so after some initial period, the trajectory is described as an equilibrium trajectory of a single, unchanging single-source single-sink instance. Theorem 3.3 can thus be directly applied. ∎
@@
In the following, $\mathcal{S}_{\theta}$, $\mathcal{Z}_{\theta}$ etc. refer to the trajectory $\ell$. We choose $\eta$ small enough that during the interval $[0,\eta]$, no new source will become active, and the set $\mathcal{Z}_{\theta}$ does not change. To guarantee this, consider the minimal slack $\tilde{s}$ (with respect to $\ell$) of an arc leaving $\mathcal{Z}_{0}$, and choose $\eta=\tilde{s}/(4\kappa)$. Then by Lipschitz continuity, any trajectory starting within distance $\tilde{s}/4$ of $l^{\circ}$ will remain within distance $\tilde{s}/2$ of $l^{\circ}$ throughout $[0,\eta]$. Since $\ell^{(i)}(0)\to l^{\circ}$, for any $i$ large enough we have that $\ell^{(i)}(\theta)$ remains within distance $\tilde{s}/2$ of $l^{\circ}$, and hence that all arcs leaving $\mathcal{Z}_{0}$ still have slack. This implies in particular that the set of active sources for $\ell^{(i)}$ is a subset of the active sources of $\ell$, for all $i$ large enough.
There may be some sources that are active for $\ell$ but inactive for $\ell^{(i)}$ at time 00. We first argue that all such sources will become active very quickly. For suppose $s_{i}$ is such a source. We have that $\beta^{(i)}_{s_{i}}(0)\leq\left\lvert V\right\rvert(\ell^{(i)}(0)-l^{\circ})$, where $\beta^{(i)}$ is defined with respect to $\ell^{(i)}$. By Lemma 7.1, this implies that $s_{i}$ becomes active by time $T^{(i)}:=(\ell^{(i)}(0)-l^{\circ})\left\lvert V\right\rvert/\kappa_{\min}$.
Thus, from time $T^{(i)}$ both $\ell^{(i)}$ and $\ell$ share the same set of active sources and the same set of nodes reachable from active sources, as long as $i$ is sufficiently large. So $\ell$ and each $\ell^{(i)}$ (for $i$ large enough) can all be viewed as equilibrium trajectories in the same single-source single-sink instance; for $\ell$ this is true for the interval $[0,\eta]$, and for $\ell^{(i)}$ on the interval $[T^{(i)},\eta]$. Further, $T^{(i)}\to 0$, meaning that $\ell^{(i)}(T^{(i)})\to\ell(T^{(i)})$. Applying Theorem 5.3 to this common instance, we deduce that $\ell^{(i)}$ converges to $\ell$ on the interval $[0,\eta)$.
Local continuity implies both uniqueness and global continuity. The argument is the same as in the single-source-single-sink setting (see Section 5, in particular proof of Theorems 5.2 and 5.3), and we do not repeat it here.
@@ 8 Conclusion
Consider a variant of the fluid queueing model where each agent is not infinitesimally small, but controls a “packet” of small but finite size. Is it true that the equilibrium of this packet model converges to the equilibrium of the fluid queueing model, as the packet size goes to 00?
@@
While we have restricted our attention to the setting of constant inflow, our uniqueness and continuity results carry over, essentially immediately, to the setting of piecewise-constant time-varying inflows. Additional technical work would be required to handle arbitrary inflow functions.
@@
Finally, a major shortcoming of our results is that they apply only to single-commodity instances. Our understanding of the multi-commodity version of the model is much less developed in general; existence of equilibria is known [CCL15]), and [SS18, Ser20] discuss some structural aspects. Can the approach taken here be developed in the multi-commodity setting? Progress here would be very exciting.
@@ Acknowledgments.
We are grateful to José Correa, Andrés Cristi, Dario Frascaria, Marcus Kaiser, Tim Oosterwijk and Philipp Warode for many interesting discussions on various topics related to equilibria in flows over time. The first author thanks Dario Frascaria in addition for useful discussions some years prior to this work on the topic of long-term behavior without a bounded inflow assumption.

### v3

@@ 7 Single-commodity instances with multiple origins and destinations
Unfortunately, we cannot handle multiple sources so easily. To see why the situation is different here, consider a setting with two sources $s_{1}$ and $s_{2}$ both connected to some vertex $v$, which is then connected to the sink $t$. Let us assume that the travel time from $s_{1}$ to $v$ is much shorter than the travel time from $s_{2}$ to $v$. In this case the particles which enter the network via $s_{2}$ at time 0 interact with particles which enter the network via $s_{1}$ at a much later time. If we were to contract $s_{1}$ and $s_{2}$ together, no particles would depart from $s_{2}$ at time 0, which is certainly incorrect.
Fortunately, the question of how to handle multiple source has already been discussed by Sering and Skutella [SS18]. They showed that Nash flows over time exist in this setting, and show that their label derivatives satisfy a modified version of the thin flow equations almost everywhere. Their formalization defines labels $\ell_{v}(\theta)$, but since there is no common source, it is not possible to identify $\theta$ as a departure time from a source, and indeed it can no longer be interpreted as a time at all. Nonetheless, $\ell_{v}(\theta)$ retains its interpretation as the earliest arrival time for a particle indexed by $\theta$, and they show that an equilibrium is described by the usual requirement that all flow is supported on $E^{\prime}_{\theta}$ (which is defined as before, via (1) and $E^{\prime}_{\theta}:=E^{\prime}_{\ell(\theta)}$). For each source $s$ and time $\xi$ there is a $\theta$ for which $\ell_{s}(\theta)=\xi$, and so this ensures that from every source, all particles traverse quickest paths to the sink. They also show that the label derivatives satisfy a variant of the thin flow equations.
We cannot quite use their results in a black box fashion, since we would like to simply apply our results for the single-source single-sink case, rather then redoing the proofs with slightly different thin flow equations.
Let $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{m}\}$ be the set of sources (and $t$ the sink, as usual), and let $u_{i}$ be the inflow rate at $s_{i}$. It will be convenient to define the following network $\hat{G}$. Its node set V(G^)V($\hat${G}) consists of $V$ along with an additional “super-source” $s$. Its arc set E(G^)E($\hat${G}) consists of $E$, along with the arcs $E_{s}:=\{ss_{i}:s_{i}\in\mathcal{S}\}$. All arcs in $E$ maintain their transit times and capacities; arc $ss_{i}$ has capacity $u_{i}$ and transit time 0 for each $s_{i}\in\mathcal{S}$. Let $u_{0}\coloneqq\sum_{i=1}^{m}u_{i}$.
We will need to consider the details of the modified thin flow equations of [SS18, Definition 3]. We will consider maximal solutions to the following, which, as $E_{\theta}^{\prime}$ is acyclic, can easily be seen to be equivalent to the modified thin flow equations of [SS18, Definition 3].
@@
$\ell^{\prime}_{s_{i}}$ $=x^{\prime}_{ss_{i}}/u_{i}$ $\text{ for all }s_{i}\in\mathcal{S}$
$\ell^{\prime}_{w}$ $\mathbin{{\leq}}\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})$ ${\text{ for all }e=vw\in E_{\theta}^{\prime}},$
$\ell^{\prime}_{w}$ $=\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})$ $\text{ for all }e=vw\in E^{\prime}_{\theta}\text{ with }x^{\prime}_{e}>0.$
@@
Call a source $s_{i}$ active at $\theta$ if there is a path in $(V,E^{\prime}_{\theta})$ from $s_{i}$ to $t$. Let $\mathcal{S}_{\theta}$ is the set of active sources at time $\theta$. Now let $\mathcal{Z}_{\theta}$ be the set of nodes reachable from the inactive sources $\mathcal{S}\setminus\mathcal{S}_{\theta}$ along a path of active arcs. Then it is easy to observe that in the modified thin flow equations, an inactive source $s_{i}$ must have $x^{\prime}_{ss_{i}}=\ell^{\prime}_{s_{i}}=0$, since there is no flow through $s_{i}$ supported on $E^{\prime}_{\theta}$. It also then follows that all nodes in $v\in\mathcal{Z}_{\theta}$ satisfy $\ell^{\prime}_{v}=0$. Let $E^{\prime\prime}_{\theta}=E^{\prime}_{\theta}\cup\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$, and let $E^{**}_{\theta}:=E^{*}_{\theta}\cup\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$. Then the following is equivalent to (10):
@@
$\ell^{\prime}_{w}$ $\mathbin{{\leq}}\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})\quad$  $\hskip-20.00003pt{\text{ for all }e=vw\in E^{\prime\prime}_{\theta}}\text{ with }w\in V\setminus\mathcal{Z}_{\theta}$
$\ell^{\prime}_{w}$ $=\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})$  $\hskip-20.00003pt\text{ for all }e=vw\in E^{\prime\prime}_{\theta}\text{ with }x^{\prime}_{e}>0{\text{ and }w\in V\setminus\mathcal{Z}_{\theta}}$
$\ell^{\prime}_{v}$ $=0$  $\hskip-20.00003pt\text{ for all }v\in\mathcal{Z}_{\theta}.$
Here, we have extended $\rho_{e}$ to $e=ss_{i}$ for $s_{i}\in\mathcal{S}_{\theta}$ in the obvious way $\rho_{e}(\ell^{\prime}_{s},x^{\prime}_{e})=x^{\prime}_{e}/u_{i}$, given that these are all considered resetting arcs. Now define the generalized network $G_{\theta}$ by first removing $\mathcal{Z}_{\theta}$ from $\hat{G}$, and then choosing the set of free arcs to be $E^{\infty}_{\theta}:=\{ss_{i}:s_{i}\in\mathcal{S}_{\theta}\}$. Note that $(E(G_{\theta}),E^{\infty}_{\theta})$ is a valid configuration, since no arc $ss_{i}$ lies on a directed cycle. Then (11) is identical to the usual thin flow equations (2) on $G_{\theta}$, along with $\ell^{\prime}_{v}=0$ for $v\in\mathcal{Z}_{\theta}$. We also fix $\ell_{s_{i}}(\theta)=0$ for all inactive sources at time $\theta$ (these labels are not meaningful); we include this as a requirement for a labeling to be feasible. As usual, we can consider equilibrium trajectories starting from any feasible labeling $l^{\circ}$.
@@
$\beta_{s_{i}}(\theta)=\min\Bigl\{\sum_{e\in P}s_{e}(\theta):P\text{ is an $s_{i}$-$t$-path}\Bigr\}.$
@@
    For all points of differentiability $\theta$ and $s_{i}\in\mathcal{S}\setminus\mathcal{S}_{\theta}$, we have $\beta_{s_{i}}^{\prime}(\theta)\leq-\kappa_{\min}$.
@@
    For any $e=vw\in\tilde{P}$, since $e$ has no queue, we have $s^{\prime}_{e}(\theta)=\ell_{v}^{\prime}(\theta)-\ell_{w}^{\prime}(\theta)$ (recall that $\theta$ is a point of differentiability). Define $\kappa_{\min}$ to be the minimum possible value of $\ell^{\prime}_{v}$ for any node $v$ and any thin flow solution $\ell^{\prime}$ in $G_{\theta}$, over all possible valid configurations (this is strictly positive by Section 2.2 and since no node of $\mathcal{Z}_{\theta}$ is in $G_{\theta}$). Then
@@
    Note that a source $s_{i}$ is active when $\beta_{s_{i}}(\theta)=0$. It follows directly that the set of active sources $\mathcal{S}_{\theta}$ only increases in $\theta$, since $\beta_{s_{i}}$ is continuous, nonnegative and monotone.
    Next, observe that for every inactive source $s_{i}$ the maximal slack on a path to the sink is upper bounded by $\sum_{e\in E}\tau_{e}$. Let $P$ be an $s_{i}$-$t$-path of minimum slack at time 0 and let $\tilde{P}$ be a maximal prefix of $P$ such that no arc of $\tilde{P}$ has a queue. Then (similarly to the above)
@@
By Section 7, we have that for all $\theta\geq\sum_{e\in E}\tau_{e}/\kappa_{\min}$ all sources are active and remain active. Thus $\mathcal{Z}_{\theta}=\emptyset$, and so after some initial period, the trajectory is described as an equilibrium trajectory of a single, unchanging single-source single-sink instance. Theorem 3.3 can thus be directly applied. ∎
@@
In the following, $\mathcal{S}_{\theta}$, $\mathcal{Z}_{\theta}$ etc. refer to the trajectory $\ell$. We choose $\eta$ small enough that during the interval $0,\eta]$, no new source will become active, and the set $\mathcal{Z}_{\theta}$ does not change. To guarantee this, consider the minimal slack $\tilde{s}$ (with respect to $\ell$) of an arc leaving $\mathcal{Z}_{0}$, and choose $\eta=\tilde{s}/(4\kappa)$. Then by Lipschitz continuity, any trajectory starting within distance $\tilde{s}/4$ of $l^{\circ}$ will remain within distance $\tilde{s}/2$ of $l^{\circ}$ throughout $0,\eta]$. Since $\ell^{(i)}(0)\to l^{\circ}$, for any $i$ large enough we have that $\ell^{(i)}(\theta)$ remains within distance $\tilde{s}/2$ of $l^{\circ}$, and hence that all arcs leaving $\mathcal{Z}_{0}$ still have slack. This implies in particular that the set of active sources for $\ell^{(i)}$ is a subset of the active sources of $\ell$, for all $i$ large enough.
There may be some sources that are active for $\ell$ but inactive for $\ell^{(i)}$ at time 0. We first argue that all such sources will become active very quickly. For suppose $s_{i}$ is such a source. We have that $\beta^{(i)}_{s_{i}}(0)\leq\left\lvert V\right\rvert(\ell^{(i)}(0)-l^{\circ})$, where $\beta^{(i)}$ is defined with respect to $\ell^{(i)}$. By Section 7, this implies that $s_{i}$ becomes active by time $T^{(i)}:=(\ell^{(i)}(0)-l^{\circ})\left\lvert V\right\rvert/\kappa_{\min}$.
Thus, from time $T^{(i)}$ both $\ell^{(i)}$ and $\ell$ share the same set of active sources and the same set of nodes reachable from active sources, as long as $i$ is sufficiently large. So $\ell$ and each $\ell^{(i)}$ (for $i$ large enough) can all be viewed as equilibrium trajectories in the same single-source single-sink instance; for $\ell$ this is true for the interval $0,\eta]$, and for $\ell^{(i)}$ on the interval $T^{(i)},\eta]$. Further, $T^{(i)}\to 0$, meaning that $\ell^{(i)}(T^{(i)})\to\ell(T^{(i)})$. Applying Theorem 3.2 to this common instance, we deduce that $\ell^{(i)}$ converges to $\ell$ on the interval $0,\eta)$.
Local continuity implies both uniqueness and global continuity. The argument is the same as in the single-source-single-sink setting (see Section 5, in particular proof of Theorems 3.1 and 3.2), and we do not repeat it here.
@@ 8 Conclusion
Consider a variant of the fluid queueing model where each agent is not infinitesimally small, but controls a “packet” of small but finite size. Is it true that the equilibrium of this packet model converges to the equilibrium of the fluid queueing model, as the packet size goes to 0?
@@
As already discussed, our results on uniqueness and continuity carry over to piecewise-constant time-varying inflows. We expect that the claims hold for arbitrary (appropriately well-behaved) inflow functions, but we do not know how to extend the arguments of this paper to show this, and we leave this as an open question.
@@
Finally, a major shortcoming of our results is that they apply only to single-commodity instances. Our understanding of the multi-commodity version of the model is much less developed in general; existence of equilibria is known [CCL15]), and [SS18, Ser20] discuss some structural aspects. Interestingly, in [Iry11] it was shown that equilibria need not be unique anymore in multi-commodity instances. It remains open whether a similar approach as taken here to show long term behaviour can be developed in the multi-commodity setting. Progress here would be very exciting.
@@ Acknowledgments.
We are grateful to José Correa, Andrés Cristi, Dario Frascaria, Marcus Kaiser, Tim Oosterwijk and Philipp Warode for many interesting discussions on various topics related to equilibria in flows over time. The first author thanks Dario Frascaria in addition for useful discussions some years prior to this work on the topic of long-term behavior without a bounded inflow assumption. Finally, we thank the anonymous referees for their detailed and constructive feedback, which led to a number of improvements.